10 km distance

1.     Based on THE REACTION BELOW – Using the amount of fuel you calculated in grams to get you from school and back, how many grams of Carbon Dioxide is produced if there is an unlimited supply of oxygen? This reaction is assuming that your car burns fuel in a complete combustion reaction… which it probably doesn’t.. but do it anyways…

2 C8H18 (g)  + 25 O2 (g)  à 16 CO2 (g)  + 18 H2(g)

 

 2.     Based on THE REACTION BELOW – Generally cars burn fuel with an unlimited supply of oxygen gas in the atmosphere. Assuming for some silly reason there was only 10.00 grams of oxygen available. Your car would stop working when the oxygen runs out, but that would be the least of your worries! Based on the amount of fuel you calculated at the start, and the 10.00g of oxygen, how many grams of water would be produced?

2 C8H18 (g)  + 17 O2 (g)  à 16 CO (g)  + 18 H2(g)

3.     Based off what happened above, how much gas in grams would be NOT used due to the lack of oxygen (solve for excess C8H18 in grams)




1
Expert's answer
2022-04-29T04:50:02-0400

1) m(C8H18)= 75g, n(C8H18)=75/114=0.67 mol, m(CO2)=n(CO2)*M(CO2)=3.35*44=147,4

2)5.96 g of water

3)72.32 g in excess